### v1

@@ Abstract
Results. Interstellar comet 3I/ATLAS has a visible spectrum consistent with that of a D-type asteroid, and has a spectral slope =14.60.2 %/1000 Å in the 5000–9000 Å range, which is similar to the one of ‘Oumuamua but redder than that of 2I/Borisov and most solar system comets. It has a conspicuous coma and its rotation period is 16.790.23 h. The heliocentric components of its Galactic velocity were km s-1 with a radiant in Sagittarius (2950500008, 19069300012). The analysis of a sample of kinematic analogs of 3I/ATLAS extracted from Gaia DR3 suggests that its parent system is part of the Galactic thin disk and includes a solar-like star with slightly sub-solar metallicity.
@@
Interstellar objects are like cosmic bottled messages that convey information about the planetary systems where they formed for anyone to analyze, connecting distant worlds meant to never interact. Any debris in the form of asteroids and comets that escapes a mature planetary system will move with values of its heliocentric Galactic velocity components similar to those of its parent star. In the absence of significant scattering due to close encounters with other stars, this kinematic signature will remain relatively unaltered for some time. At least two comets — C/1980 E1 (Bowell) and C/2024 L5 (ATLAS) — have been recently ejected from the solar system after experiencing planetary encounters at close range (de la Fuente Marcos et al., 2024). In contrast, we had two visitors from beyond the solar system — 1I/2017 U1 (‘Oumuamua) and 2I/Borisov — and they left more questions raised than answered (see, e.g., Williams 2017; Meech et al. 2017; Micheli et al. 2018; de León et al. 2019, 2020; Deam et al. 2025). The study of more of these visitors from beyond the solar system will only broaden our understanding of this subject.
Object A11pl3Z was initially reported on July 1, 2025, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Rio Hurtado in Chile. It was announced as the third interstellar object on July 2 (Collaboration et al., 2025b) and designated 3I/ATLAS = C/2025 N1 (ATLAS). Follow-up and precovery observations (Ye et al., 2025b, a; Valdes et al., 2025; Collaboration et al., 2025a) revealed a highly eccentric, hyperbolic orbit (eccentricity 6). Here, we investigate observationally the spectral properties of 3I/ATLAS in the visible region using low-resolution spectroscopy, its cometary activity using high SNR images, and its rotational state from photometric observations. Our goal is to place 3I/ATLAS within the context of both known interstellar objects and solar system comets. In addition, we use its orbital solution to explore numerically its dynamical evolution. This Letter is organized as follows. In Sect. 2, we analyze the observed properties of 3I/ATLAS, comparing with those of known interstellar objects and solar system comets. Our results are discussed in Sect. 3 and our conclusions are summarized in Sect. 4. Observations, data reduction, spectral extraction, and other technical details and input data are included as appendices.
@@ 2.1 Cometary activity
We generated a deep, stacked image by co-adding 227 exposures (a total of 3.15 h) from the TTT telescope obtained on the night of 2 July, that is shown in Fig. 1. Overlaid contour delineates the coma boundary at the level above the background sky. At this threshold, the surface brightness of the edge is mag arcsec-2 in the AB magnitude system. Details of the reduction and photometric calibration are given in Appendix B. The comet shows a well‐defined coma, extending over a projected area of . The major axis of this ellipse-like region is oriented at a phase angle of . At the time of observation, the geocentric distance was 3.44 au, so the physical dimensions of approximately 26 400 km 24 700 km.
@@ 2.2 Spectral properties
We studied the presence of typical cometary emission lines/bands in cometary comas (see. e.g.. Cochran et al. 2012). To do this, we proceeded as in de León et al. (2020), we obtained and analyzed the continuum-subtracted spectrum of 3I/ATLAS (see Fig. 5). We first noticed that no cometary gas emission was detected as also reported by Alvarez-Candal et al. (2025) and Opitom et al. (2025), thus the observed coma is mainly made of dust. Then we derived an upper limit for the most prominent gas emission band of comets in this spectral range, the CN (0–0) at 3880 Å, and obtained mol s-1 using the method described in Appendix C. Our results are consistent with the value reported in Alvarez-Candal et al. (2025). Ours is larger because the of their spectrum in the CN band spectral range is larger as a result of their much higher spectral resolution. Notice that measured production rates mol s-1 are unusual (A’Hearn et al., 1995) so 3I/ATLAS looks like a normal comet at this solar distance.
On the other hand, the normalized reflectance spectrum of interstellar comet 3I/ATLAS is compatible with a D-type classification, following DeMeo et al. (2009), and similar to the red spectra of the trans-Neptunian objects and Centaurs. We computed its reflectivity gradient by using a simple linear fit to its reflectance spectrum, with the same wavelength intervals and normalization wavelength as in Opitom et al. (2025) for comparison purposes: (18.50.5)%/1000 Å in the 5000-7000 Å range, (11.60.4)%/1000 Å in the 7000-9000 Å range, and (14.60.2)%/1000 Å in the 5000-9000 Å range.
@@ 2.4 Dynamics
The kinematic properties of interstellar interlopers are expected to preserve for the most part the kinematic signatures of their host stars. Analyzing the pre-encounter trajectory of 3I/ATLAS might shed some light on its true origin. We performed integrations backward in time of 1000 control orbits of 3I/ATLAS generated by the Monte Carlo using the Covariance Matrix (MCCM) method described by de la Fuente Marcos & de la Fuente Marcos (2015) and modified here to work with hyperbolic orbits. A statistical analysis (median and 16th and 84th percentiles) of the results indicates that at pc from the Sun and 3104 yr into the past, 3I/ATLAS was moving inwards, at 58.000.02 km s-1 — 1I/2017 U1 (‘Oumuamua) had km s-1, e.g. Mamajek (2017) and C/2019 Q4 (Borisov) had km s-1 (de León et al., 2020) — and projected towards (radiant or antapex) , (2950500008, ) in the constellation of Sagittarius with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity were km s-1. These values were computed as described by de la Fuente Marcos & de la Fuente Marcos (2019). The components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS are (62.40.8, 7.20.5, 26.20.4) km s-1. We have searched Gaia third data release (DR3, Gaia Collaboration et al. 2016, 2023) looking for kinematic analogues of 3I/ATLAS towards the region of the radiant. Figure 4 shows 3I/ATLAS in its Galactic kinematic context.
Answering the question of where 3I/ATLAS is headed next after its fly-by with the Sun requires the analysis of a similar set of direct -body simulations (also with MCCM), but forward in time. At pc from the Sun and 3104 yr into the future, this object will be receding from us at 58.010.02 km s-1 towards (apex) , (9526003, ) in the constellation of Gemini with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity will be km s-1.
@@ 3 Discussion
Our spectroscopic results are in agreement with the ones reported by Opitom et al. (2025), and are of the order of those reported in Alvarez-Candal et al. (2025). In addition, Seligman et al. (2025) reported a slope of 18%/1000 Å from their ’,’,’,’ colors and (17.10.2)%/1000 Å from their spectrum, consistent with the values provided in this work. The spectral slope of 3I/ATLAS is redder than the nominal values adopted by Jewitt & Seligman (2023) for 1I/’Oumuamua (155 %/1000 Å) and 2I/Borisov (121 %/1000 Å).
Considering Nissen (2004), fig. 1.3, the components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS place it in the thin disk as the value of is 67.68 km s-1 and =7.2 km s-1. The Toomre diagram indicates that the thin disk is inside 85 km s-1 with the thick disk in the interval (85, 180) km s-1 and the halo outside 180 km s-1. In addition to this kinematic constraint, the analysis of a sample of kinematic analogues of 3I/ATLAS from Gaia DR3 gives a value of [Fe/H]=0.040.14 (see Appendix F), which is somewhat sub-solar, but far from 1.2[Fe/H]0.4 in the thick disk (see, e.g., Nissen 2004). An origin in the thick disk is favored by Hopkins et al. (2025).
@@ 4 Conclusions
3I/ATLAS has a visible reflectance spectrum somewhat consistent with that of Solar System comets and D-type asteroidd, with a spectral slope of (18.50.5)%/1000 Å in the 5000–7000Å range, (11.60.4)%/1000 Å in the 7000–9000Å range, and (14.60.2)%/1000 Å in the 5000–9000 Å range.
@@
No signatures of gas emission is observed in 3I/ATLAS visible spectrum, which is usual in solar system comets observed at au. We obtained an upper limit for the CN production rate mol s-1.
@@
The analysis of an extensive set of -body simulations indicate that 3I/ATLAS came from interstellar space and it will return back to it after reaching its perihelion on 2025 October 29. Its kinematic signature is consistent with an origin in the Galactic thin disk.
@@ Acknowledgements.
RdlFM acknowledges funding from the “ENIMUS” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement ID 101097905). This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. This article is based on observations made in the Two-meter Twin Telescope (TTT111http://ttt.iac.es) sited at the Teide Observatory of the Instituto de Astrofísica de Canarias (IAC), that Light Bridges operates in Tenerife, Canary Islands (Spain). The Observing Time Rights (DTO) used for this research were consumed in the PEI “PLANETIX25”. This research used storage and computing capacity in ASTRO POC’s EDGE computing center at Tenerife under the form of Indefeasible Computer Rights (ICR). The ICR were consumed in the PEI ”PLANETIX25” with the collaboration of Hewlett Packard Enterprise and VAST DATA. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Collaboration et al. (2025a) Collaboration, Z. T. F., Ye, Q. Z., Nastasi, A., et al. 2025a, Minor Planet Electronic Circulars, 2025-N51
- Collaboration et al. (2025b) Collaboration, Z. T. F., Ye, Q. Z., Robinson, J., et al. 2025b, Minor Planet Electronic Circulars, 2025-N12
@@ Appendix A Observations and Data Reduction
Observations of interstellar comet 3I/ATLAS were performed soon after its discovery and confirmation of interstellar nature, using the telescope facilities located at the Canarian Observatories (OCAN), managed by the Instituto de Astrofísica de Canarias (Spain). Observational circumstances are described in the following sections and summarized in Table 1.
@@ A.1 Spectroscopy
Spectroscopic observations of 3I/ATLAS in the near-UV and visible wavelengths (3500–9200 Å) were obtained using OSIRIS+ camera-spectrograph (Cepa et al., 2000; Cepa, 2010) at the 10.4m Gran Telescopio Canarias (GTC), located at the El Roque de Los Muchachos Observatory (La Palma, Spain). The OSIRIS+ instrument was upgraded in January 2023 to have enhanced sensitivity in the blue. It is equipped with 2k4k pixel detector that provides a 7.87.8 arcmin2. We used the R300B (3500–7500 Å, dispersion of 4.96 Å pixel-1 for a 0.6” slit) and R300R (4800–9200 Å, dispersion of 7.74 Å pixel-1 for a 0.6” slit) grisms and a 1.5” slit width oriented in the parallactic angle. The telescope tracking was at the comet’s proper motion. Two spectrophotometric standard stars (Feige 110 and Ross 640) were observed to flux calibrate the spectra. In the same way, to obtain the relative reflectance spectrum of the comet, we also observed two solar analog stars — SA110-361 and SA112-1333 — from the Landolt catalog (Landolt, 1992).
Data reduction included bias and flat-field correction, sky background subtraction and one-dimensional aperture extraction, corresponding to the pixel where the peak intensity of the cometary profile decayed to its 10% value. Wavelength calibration was then applied to the one-dimensional spectra using Xe+Ne+HgAr lamps. This procedure was applied to the spectra of the comet and the stars (both spectrophotometric and solar analogs). We divided the individual spectra of the comet by the spectra of the solar analogs, and the resulting ratio were averaged to get the final reflectance spectra. To create the final, composite reflectance spectrum shown in Fig. 2, we joined the R300B and R300R using the common wavelength interval (6000–7000 Å), and normalized the result to unity at 5500 Å.

### v2

@@ Abstract
Results. Interstellar comet 3I/ATLAS has a visible spectrum slightly redder than those of D-type asteroids, and has a spectral slope =14.60.2 %/1000 Å in the 5000–9000 Å range, which is similar to the one of ‘Oumuamua but redder than that of 2I/Borisov and most solar system comets. It has a conspicuous coma and its rotation period is 16.790.23 h. The heliocentric components of its Galactic velocity were km s-1 with a radiant in Sagittarius ). The analysis of a sample of kinematic analogs of 3I/ATLAS extracted from Gaia DR3 suggests that its parent system is part of the Galactic thin disk and includes a solar-like star with slightly sub-solar metallicity.
@@
Interstellar objects are like cosmic messages in a bottle that convey information about the planetary systems where they formed for anyone to analyze, connecting distant worlds meant to never interact. Any debris in the form of asteroids and comets that escapes a mature planetary system will move with values of its heliocentric Galactic velocity components similar to those of its parent star. In the absence of significant scattering due to close encounters with other stars, this kinematic signature will remain relatively unaltered for some time. At least two comets — C/1980 E1 (Bowell) and C/2024 L5 (ATLAS) — have been recently ejected from the solar system after experiencing planetary encounters at close range (de la Fuente Marcos et al., 2024). In contrast, we had two visitors from beyond the solar system — 1I/2017 U1 (‘Oumuamua) and 2I/Borisov — and they left more questions raised than answered (see, e.g., Williams 2017; Meech et al. 2017; Micheli et al. 2018; de León et al. 2019, 2020; Deam et al. 2025). The study of more of these visitors from beyond the solar system will only broaden our understanding of this subject.
Object A11pl3Z was initially reported on July 1, 2025, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Rio Hurtado in Chile. It was announced as the third interstellar object on July 2 (Denneau et al., 2025) and designated 3I/ATLAS = C/2025 N1 (ATLAS). Follow-up and precovery observations (Ye et al., 2025b, a; Valdes et al., 2025; Collaboration et al., 2025; Bellini et al., 2025; Chernyshov et al., 2025; Ye et al., 2025c) revealed a highly eccentric, hyperbolic orbit (eccentricity 6). Here, we investigate observationally the spectral properties of 3I/ATLAS in the visible region using low-resolution spectroscopy, its cometary activity using high SNR images, and its rotational state from photometric observations. Our goal is to place 3I/ATLAS within the context of both known interstellar objects and solar system comets. In addition, we use its orbital solution to explore numerically its dynamical evolution. This Letter is organized as follows. In Sect. 2, we analyze the observed properties of 3I/ATLAS, comparing with those of known interstellar objects and solar system comets. Our results are discussed in Sect. 3 and our conclusions are summarized in Sect. 4. Observations, data reduction, spectral extraction, and other technical details and input data are included as appendices.
@@ 2.1 Cometary activity
We generated a deep, stacked image by co-adding 227 exposures (a total of 3.15 h) from the TTT telescope obtained on the night of 2 July, that is shown in Fig. 1. Overlaid contour delineates the coma boundary at the level above the background sky. At this threshold, the surface brightness of the edge is mag arcsec-2 in the AB magnitude system. Details of the reduction and photometric calibration are given in Appendix B. The comet shows a well‐defined coma, extending over a projected area of . The major axis of this ellipse-like region is oriented at a phase angle of . At the time of observation, the geocentric distance was 3.44 au, so the physical dimensions were approximately 26 400 km 24 700 km.
Using a circular aperture of radius (10 000 km) from the centroid, we measured the integrated flux and converted it into the dust‑production proxy following the formulation of A’Hearn et al. (1984). After correcting this quantity for the phase angle with the Halley–Marcus composite dust phase function (Schleicher & Bair, 2011), we obtain a phase‑normalised ‑band value of cm. This is slightly larger than the cm reported by Bolin et al. (2025) for a comparable aperture.
@@ 2.2 Spectral properties
We studied the presence of typical cometary emission lines/bands in cometary comas (see. e.g.. Cochran et al. 2012). To do this, we proceeded as in de León et al. (2020), we obtained and analyzed the continuum-subtracted spectrum of 3I/ATLAS (see Fig. 4). We first noticed that no cometary gas emission was detected as also reported by Alvarez-Candal et al. (2025) and Opitom et al. (2025), thus the observed coma is mainly made of dust. Then we derived an upper limit for the most prominent gas emission band of comets in this spectral range, the CN (0–0) at 3880 Å, and obtained mol s-1 using the method described in Appendix C. Our results are consistent with the value reported in Alvarez-Candal et al. (2025). Ours is larger because the of their spectrum in the CN band spectral range is better as a result of their much higher spectral resolution. Notice that measured CN production rates at large heliocentric distances are unusual (A’Hearn et al., 1995) so 3I/ATLAS looks like a normal comet.
On the other hand, the normalized reflectance spectrum of interstellar comet 3I/ATLAS, shown in Fig. 2 is slightly redder than D-types, following DeMeo et al. (2009), and similar to the red spectra of the trans-Neptunian objects and Centaurs. For comparison purposes, we also show in the plot the normalized spectra of 1I/’Oumuamua (Fitzsimmons et al., 2018) and 2I/Borisov (de León et al., 2020). We computed the reflectivity gradient by using a simple linear fit to the reflectance spectrum of 3I/ATLAS in the wavelength range 3800-9200 Å, obtaining a value of ’=(18.30.9)%/1000 Å. Spectral slope error corresponds to the standard deviation of the mean of a total of 100 iterations, randomly removing 10 points during each fit. We are not considering here the systematic error of the observations associated to the use of two different solar analogs (see Appendix A.1). Using the same wavelength intervals and normalization wavelength (6000 Å) as in Opitom et al. (2025), we obtain: (18.50.5)%/1000 Å in the 5000-7000 Å range, (11.60.4)%/1000 Å in the 7000-9000 Å range, and (14.60.2)%/1000 Å in the 5000-9000 Å range.
@@ 2.4 Dynamics
The kinematic properties of interstellar interlopers are expected to preserve, for the most part, the kinematic signatures of their host stars. Analyzing the pre-encounter trajectory of 3I/ATLAS might shed some light on its true origin. We performed integrations backward in time of 1000 control orbits of 3I/ATLAS generated by the Monte Carlo using the Covariance Matrix (MCCM) method described by de la Fuente Marcos & de la Fuente Marcos (2015) and modified here to work with hyperbolic orbits. A statistical analysis (median and 16th and 84th percentiles) of the results indicates that at 1.78190.0003 pc from the Sun and 3104 yr into the past, 3I/ATLAS was moving inwards, at 57.980 km s-1 — 1I/2017 U1 (‘Oumuamua) had km s-1, e.g. Mamajek (2017) and C/2019 Q4 (Borisov) had km s-1 (de León et al., 2020) — and projected towards (radiant or antapex) , (295043, ) in the constellation of Sagittarius with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity were km s-1. These values were computed as described by de la Fuente Marcos & de la Fuente Marcos (2019). The components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS are (62.30.8, 7.20.5, 26.20.4) km s-1. We have searched Gaia third data release (DR3, Gaia Collaboration et al. 2016, 2023) looking for kinematic analogues of 3I/ATLAS towards the region of the radiant. Figure 6 shows 3I/ATLAS in its Galactic kinematic context.
@@ 3 Discussion
Our spectroscopic results are comparable with those reported by Opitom et al. (2025) and Alvarez-Candal et al. (2025), although we obtained a slightly bluer spectral slope in the 7000-9000 Å. Interestingly, Belyakov et al. (2025) also found a more neutral slope for their spectrum in the range 7000–10000 Å, as they did in Kareta et al. (2025) at longer wavelengths. Seligman et al. (2025) reported a slope of 18%/1000 Å from their ’,’,’,’ colors and (17.10.2)%/1000 Å from their spectrum, consistent with the spectral gradient provided in this work in the 3800-9200 Å range. The spectral slope of 3I/ATLAS is redder than the nominal values adopted by Jewitt & Seligman (2023) for 1I/’Oumuamua (155 %/1000 Å) and 2I/Borisov (121 %/1000 Å).
Considering Nissen (2004), fig. 1.3, the components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS place it in the thin disk as the value of is 67.58 km s-1 and =7.2 km s-1. The Toomre diagram indicates that the thin disk is inside 85 km s-1 with the thick disk in the interval (85, 180) km s-1 and the halo outside 180 km s-1 (see fig. 3 in Bensby et al. 2003). In addition to this kinematic constraint, the analysis of a sample of kinematic analogues of 3I/ATLAS from Gaia DR3 gives a value of [Fe/H]=0.040.14 (see Appendix F), which is somewhat subsolar, but far from 1.2[Fe/H]0.4 in the thick disk (see, e.g. Nissen 2004). An origin in the thick disk is favored by Hopkins et al. (2025). Taylor & Seligman (2025) argue for an age range 3–11 Gyr and a value of [Fe/H], both compatible with thin disk membership.
@@ 4 Conclusions
3I/ATLAS has a visible reflectance spectrum slightly redder than the spectra of 1I/’Oumuamua and 2I/Borisov, but consistent with that of Solar System comets, TNOs and Centaurs. Its overall spectral gradient in the 3800-9200 Å, ’=(18.30.9)%/1000 Å, is in good agreement with the findings by other authors, although somewhat bluer in the 7000-5000 Å range (’=11.60.4)%/1000 Å), suggesting a more neutral slope to longer wavelengths.
@@
No signature of gas emission is observed in 3I/ATLAS visible spectrum, which is normal in solar system comets observed at au. We obtained an upper limit for the CN production rate mol s-1.
@@
The analysis of an extensive set of -body simulations indicates that 3I/ATLAS came from interstellar space and it will return back to it after reaching its perihelion on 2025 October 29. Its kinematic signature is consistent with an origin in the Galactic thin disk.
@@ Acknowledgements.
We thank the anonymous referee for a prompt, constructive, detailed, and actionable report. RdlFM acknowledges funding from the “ENIMUS” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement ID 101097905). This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. This article is based on observations made in the Two-meter Twin Telescope (TTT111http://ttt.iac.es) sited at the Teide Observatory of the Instituto de Astrofísica de Canarias (IAC), that Light Bridges operates in Tenerife, Canary Islands (Spain). The observation time rights (DTO) used for this research were consumed in the PEI ’PLANETIX25’. This research used storage and computing capacity in ASTRO POC’s EDGE computing center at Tenerife under the form of Indefeasible Computer Rights (ICR). The ICR were consumed in the PEI ”PLANETIX25” with the collaboration of Hewlett Packard Enterprise and VAST DATA. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- A’Hearn et al. (1984) A’Hearn, M. F., Schleicher, D. G., Millis, R. L., Feldman, P. D., & Thompson, D. T. 1984, AJ, 89, 579
@@
- Bellini et al. (2025) Bellini, F., Aletti, A., Buzzi, L., et al. 2025, Minor Planet Electronic Circulars, 2025-N89
- Belyakov et al. (2025) Belyakov, M., Fremling, C., Graham, M. J., et al. 2025, arXiv e-prints, arXiv:2507.11720
- Bensby et al. (2003) Bensby, T., Feltzing, S., & Lundström, I. 2003, A&A, 410, 527
@@
- Bolin et al. (2025) Bolin, B. T., Belyakov, M., Fremling, C., et al. 2025, Monthly Notices of the Royal Astronomical Society: Letters, slaf078
@@
- Chernyshov et al. (2025) Chernyshov, A., Nazarov, S., Robinson, J., et al. 2025, Minor Planet Electronic Circulars, 2025-N102
@@
- Collaboration et al. (2025) Collaboration, Z. T. F., Ye, Q. Z., Nastasi, A., et al. 2025, Minor Planet Electronic Circulars, 2025-N51
@@
- Denneau et al. (2025) Denneau, L., Siverd, R., Tonry, J., et al. 2025, Minor Planet Electronic Circulars, 2025-N12
- Fitzsimmons et al. (2018) Fitzsimmons, A., Snodgrass, C., Rozitis, B., et al. 2018, Nature Astronomy, 2, 133
@@
- Kareta et al. (2025) Kareta, T., Champagne, C., McClure, L., et al. 2025, arXiv e-prints, arXiv:2507.12234
@@
- Taylor & Seligman (2025) Taylor, A. G. & Seligman, D. Z. 2025, arXiv e-prints, arXiv:2507.08111
@@
- Ye et al. (2025c) Ye, Q. Z., Collaboration, Z. T. F., Serrau, M., et al. 2025c, Minor Planet Electronic Circulars, 2025-O05
@@ Appendix A Observations and Data Reduction
Observations of interstellar comet 3I/ATLAS were carried out soon after its discovery and confirmation of interstellar nature, using the telescope facilities located at the Canarian Observatories (OCAN), managed by the Instituto de Astrofísica de Canarias (Spain). Observational circumstances are described in the following sections and summarized in Table 1.
@@ A.1 Spectroscopy
Spectroscopic observations of 3I/ATLAS in the near-UV and visible wavelengths (3500–9200 Å) were obtained using OSIRIS+ camera-spectrograph (Cepa et al. 2000; Cepa 2010) at the 10.4m Gran Telescopio Canarias (GTC), located at the El Roque de Los Muchachos Observatory (La Palma, Spain). The OSIRIS+ instrument was upgraded in January 2023 to have enhanced sensitivity in the blue, down to 3700-3800 Å. It is equipped with 2k4k pixel detector that provides a 7.87.8 arcmin2. We used the R300B (3600–7500 Å, dispersion of 4.96 Å pixel-1 for a 0.6” slit) and R300R (4800–9200 Å, dispersion of 7.74 Å pixel-1 for a 0.6” slit) grisms and a 1.5” slit width oriented in the parallactic angle. The telescope tracking was at the comet’s proper motion. Two spectrophotometric standard stars (Feige 110 and Ross 640) were observed to flux calibrate the spectra. In the same way, to obtain the relative reflectance spectrum of the comet, we also observed two solar analog stars — SA110-361 and SA112-1333 — from the Landolt catalog (Landolt 1992).
Data reduction included bias and flat-field correction, sky background subtraction, and one-dimensional aperture extraction, corresponding to the pixel where the peak intensity of the cometary profile decayed to its 10% value. Wavelength calibration was then applied to the one-dimensional spectra using Xe+Ne+HgAr lamps. This procedure was applied to the spectra of the comet and the stars (both spectrophotometric and solar analogs). We divided the individual spectra of the comet by the spectra of the solar analogs, and the resulting ratio were averaged to get the final reflectance spectra. To create the final composite reflectance spectrum shown in Fig. 2, we put together the R300B and R300R data using the common wavelength interval (6000–7000 Å), and normalized the result to unity at 5500 Å. The light shaded areas in the figure correspond to the standard deviation of the average, which introduces an error in the calculation of the spectral gradient ’ smaller than 1%/1000 Å.